# Providence, USA

### v1

@@ 5 Combined Constraints on EW-portal Benchmark Models
With the model-independent bounds obtained in the previous sections, given a specific model where one can calculate the branching ratios of $Z$ and $h$ decays to the dark sector, the production cross sections of $Z^{\prime}$ and $\phi$, and the effective ALP decay constant $f_{a}$ of the dark pions. The reaches in the model parameter space from various experiments can be compared and combined directly. In this section, we will illustrate this with a couple of EW-portal benchmarks discussed in Refs. Cheng:2019yai ; Cheng:2024hvq . The first one is a $Z$-portal model where light dark quarks acquire small couplings to $Z$ by mixing with heavy EW doublet fermions through the Higgs VEV, as shown in the middle panel of Fig. 1. In the second benchmark, the couplings arise from the mixing between $Z$ and a dark $Z^{\prime}$, which couple to the light dark quarks. The Higgs-initiated dark showers are also present in both models, and the scalar $\phi$ can exist in the second model, but their contributions are subleading. The lifetimes of dark pions are typically too long for collider experiments in pure Higgs-portal models unless additional ingredients are added to enhance the dark pion decays. Therefore, we focus on $Z$- and $Z^{\prime}$-initiated dark showers together with FCNC $B$ decays. However, we emphasize that they only serve as examples but do not necessarily represent general results due to the large parameter space to be explored. For concreteness, we take the number of dark colors $N_{d}=3$ and the dark flavor number $N=2$, with $N^{2}-1=3$ light dark pions in both cases.
For the heavy doublet fermion mixing model Cheng:2021kjg , the light dark quarks $\psi_{L}$, $\psi_{R}$ couple to some heavy EW doublets $Q_{L}$, $Q_{R}$,
@@
The FCNC production of the dark pions mainly depends on the effective ALP decay constant $f_{a}$, so it is insensitive to model details (except for some logarithmic dependence on the UV completion scale) as long as it gives the same $f_{a}$. Therefore, it provides an independent probe for the dark shower search. The relative strengths in experimental bounds between the dark shower and FCNC productions can only be evaluated within specific models, which is the task of this section.
Figure 19: Model-dependent exclusion limits from the dimuon DV scouting strategy at (HL-)LHC with various channels. LEFT: Exclusion limits for a fermion doublet model. RIGHT: Exclusion limits for a $Z^{\prime}$ model with $m_{Z^{\prime}}$=20 GeV.
@@
Constraints from Auxiliary Detectors Following a similar procedure, the model-dependent limits for both benchmark models at (HL-)LHC auxiliary detectors are plotted in Fig. 21. The dark shower signals are accessible at MATHUSLA, ANUBIS, and Codex-b. As they are sensitive to dark pions with $c\tau(\hat{\pi})\gtrsim 10$ m with low backgrounds, the reach of $f_{a}$ can get close to or even exceed $10^{2}$ PeV. Meanwhile, all auxiliary detectors considered can probe the FCNC decay signals, though for $f_{\hat{\pi}}=1$ GeV, they are weaker than the dark shower reaches. Nevertheless, they can still reach $f_{a}$ values comparable to the dark shower limits in other detectors thanks to the detectors’ reach in the long-lifetime regime.
@@ 6 Summary and Discussion
The details of the dark showers produced from the EW portals mostly depend only on the parameters of the dark hadrons, but not the mediators. This allows us to take a model-independent approach in studying the reaches of the collider searches. The rate of producing observable displaced (dimuon) events is governed by the exotic decay branching ratios of $Z$, $h$ into dark quarks, or the mixing parameters between the dark $Z^{\prime}$, $\phi$ with $Z$, $h$ (and the branching ratio of dark pion decaying into dimuon). The experimental reaches of the rate depend on the dark pion masses and their lifetimes. For each experiment, we can project the reach limits on the production rate vs. decay length plane for a given dark pions mass. These limits can be easily applied to obtain constraints on model parameters for any specific models that produce EW scale dark showers. Consequently, it is a valuable exercise to provide such model-independent limits. In addition to obtaining the current limits by recasting the current searches, we also make projections on future reaches for CMS scouting searches and LHCb, as well as other proposed future detectors or experiments, based on some conservative assumptions. The projections could be further enhanced if more improvements in experimental searches can be realized or better strategies can be employed. For example, the data parking approach retains the full event information and could be even more powerful than data scouting when searching for dark shower events. However, there is no current search available for us to make realistic projections.
@@
The light dark hadrons can also be produced in FCNC $B$ decays with similar search strategies as the dark shower searches. The experimental reaches can also be studied in a model-independent way. The bounds are put on the branching ratio of FCNC $B$ decay into dark pions vs. dark pion lifetime plane for a given dark pion mass. In addition to LHC experiments discussed in the dark shower search, the FCNC decays can also be searched at fixed-target or beam-dump experiments, the $B$-factory Belle II, and the LHC forward experiments such as the FASER. The projections of their reaches are also obtained. The branching ratio of the FCNC meson decay into a dark pion mostly just depends on the effective ALP decay constant $f_{a}$, plus a mild logarithmic dependence of the UV cutoff of the underlying theory. It provides independent information from the dark shower events, which depend on their productions from the EW bosons, and therefore serves as a complementary test.
@@
The relative strengths of various experiments are model dependent. To compare them, specific models are required. The model-independent bounds can then be translated into bounds on the model parameters. We illustrate this with two benchmark models with dark showers generated dominantly from $Z$ and $Z^{\prime}$. Reach limits from different experiments can be cast in the same parameter space for comparison. We found that most experiments can probe regions beyond the current bounds allowed by the EWPT. For $f_{\hat{\pi}}=1$ GeV, the reaches of the dark shower searches are typically somewhat stronger than the FCNC decay searches, although for larger $f_{\hat{\pi}}$ the trend could be reversed. CMS scouting search and LHCb generally give comparable limits, with LHCb winning in the case of a light $Z^{\prime}$. For longer decay lengths, the auxiliary detectors, in particular, MATHUSLA, will have the most powerful reach thanks to its location and large volume. Of course, these two benchmark models only represent a small set of the large possible model space, and the comparison of different experiments within these models should not be taken as general conclusions. Nevertheless, they demonstrate how the model-independent results obtained in this paper can be applied to specific models.
@@ Appendix A Dimuon Vertex Trigger Efficiencies at HL-LHC
Appendix B Benchmarks for Model-Dependent Projections

### v2

@@ 5 Combined Constraints on EW-portal Benchmark Models
With the general bounds obtained in the previous sections, one can derive model-specific bounds on the model parameters. Given a specific model, one can calculate the branching ratios of $Z$ and $h$ decays to the dark sector, the production cross sections of $Z^{\prime}$ and $\phi$, the effective ALP decay constant $f_{a}$ of the dark pions, and the FCNC $B$ decay rates. The reaches in the model parameter space from various experiments can then be compared and combined directly. In this section, we will illustrate this with a couple of EW-portal benchmarks discussed in Refs. Cheng:2019yai ; Cheng:2024hvq . The first one is a $Z$-portal model where light dark quarks acquire small couplings to $Z$ by mixing with heavy EW doublet fermions through the Higgs VEV, as shown in the middle panel of Fig. 1. In the second benchmark, the couplings arise from the mixing between $Z$ and a dark $Z^{\prime}$, which couple to the light dark quarks. The Higgs-initiated dark showers are also present in both models, and the scalar $\phi$ can exist in the second model, but their contributions are subleading. The lifetimes of dark pions are typically too long for collider experiments in pure Higgs-portal models unless additional ingredients are added to enhance the dark pion decays. Therefore, we focus on $Z$- and $Z^{\prime}$-initiated dark showers together with FCNC $B$ decays. However, we emphasize that they only serve as examples but do not necessarily represent general results due to the large parameter space to be explored. For concreteness, we take the number of dark colors $N_{d}=3$ and the dark flavor number $N=2$, with $N^{2}-1=3$ light dark pions in both cases. For the heavy doublet fermion mixing model Cheng:2021kjg , the light dark quarks $\psi_{L}$, $\psi_{R}$ couple to some heavy EW doublets $Q_{L}$, $Q_{R}$,
@@
The FCNC production of the dark pions mainly depends on the effective ALP decay constant $f_{a}$, so it is insensitive to model details (except for some logarithmic dependence on the UV completion scale) as long as it gives the same $f_{a}$. Therefore, it provides an independent probe for the dark shower search. Since the search strategy is based on single DV without isolation requirements, the SM decay products of $B$ have no direct impact on the analysis. It is thus reasonable to assume the signal efficiency of the inclusive $B\to\hat{\pi}+X_{s}$ decays will be similar to the exclusive $B\to\hat{\pi}+K^{(\ast)}$ decay ones. According to Eq. (10), the inclusive process has a decay rate $\sim$ 7 times larger than the exclusive ones discussed. To better demonstrate the potential of the flavor probes, we will plot the inclusive $B$ decay constraints based on the assumption above. The relative strengths in experimental bounds between the dark shower and FCNC productions can only be evaluated within specific models, which is the task of this section.
Figure 19: Model-specific 95% C.L. exclusion limits from the dimuon DV scouting strategy at (HL-)LHC with various channels. LEFT: Exclusion limits for a fermion doublet model. RIGHT: Exclusion limits for a $Z^{\prime}$ model with $m_{Z^{\prime}}$=20 GeV.
@@
Constraints from Auxiliary Detectors Following a similar procedure, the model-specific limits for both benchmark models at (HL-)LHC auxiliary detectors are plotted in Fig. 21. The dark shower signals are accessible at MATHUSLA, ANUBIS, and Codex-b. As they are sensitive to dark pions with $c\tau(\hat{\pi})\gtrsim 10$ m with low backgrounds, the reach of $f_{a}$ can get close to or even exceed $10^{2}$ PeV. Meanwhile, all auxiliary detectors considered can probe the FCNC decay signals, though for $f_{\hat{\pi}}=1$ GeV, they are weaker than the dark shower reaches. Nevertheless, they can still reach $f_{a}$ values comparable to the dark shower limits in other detectors thanks to the detectors’ reach in the long-lifetime regime.
@@ 6 Summary and Discussion
The details of the dark showers produced from the EW portals mostly depend only on the parameters of the dark hadrons and dark QCD, but not the mediators. This allows us to study their collider reaches in a simplified approach. The rate of producing observable displaced (dimuon) events is governed by the exotic decay branching ratios of $Z$, $h$ into dark quarks, or the mixing parameters between the dark $Z^{\prime}$, $\phi$ with $Z$, $h$ (and the branching ratio of dark pion decaying into dimuon). The experimental reaches of the rate depend on the dark pion masses and their lifetimes. For each experiment, we can project the reach limits on the production rate vs. decay length plane for a given dark pions mass. These limits can be easily applied to obtain constraints on model parameters for any specific models that produce EW scale dark showers. Consequently, it is a valuable exercise to provide such limits. In addition to obtaining the current limits by recasting the current searches, we also make projections on future reaches for CMS scouting searches and LHCb, as well as other proposed future detectors or experiments, based on some conservative assumptions. The projections could be further enhanced if more improvements in experimental searches can be realized or better strategies can be employed. For example, the data parking approach retains the full event information and could be even more powerful than data scouting when searching for dark shower events. However, there is no current search available for us to make realistic projections.
@@
The light dark hadrons can also be produced in FCNC $B$ decays with similar search strategies as the dark shower searches. The experimental reaches can also be studied. The bounds are put on the branching ratio of FCNC $B$ decay into dark pions vs. dark pion lifetime plane for a given dark pion mass. In addition to LHC experiments discussed in the dark shower search, the FCNC decays can also be searched at fixed-target or beam-dump experiments, the $B$-factory Belle II, and the LHC forward experiments such as the FASER. The projections of their reaches are also obtained. The branching ratio of the FCNC meson decay into a dark pion mostly just depends on the effective ALP decay constant $f_{a}$, plus a mild logarithmic dependence of the UV cutoff of the underlying theory. It provides independent information from the dark shower events, which depend on their productions from the EW bosons, and therefore serves as a complementary test.
@@
The relative strengths of various experiments are model dependent. To compare them, specific models are required. The bounds can then be translated into bounds on the model parameters. We illustrate this with two benchmark models with dark showers generated dominantly from $Z$ and $Z^{\prime}$. Reach limits from different experiments can be cast in the same parameter space for comparison. We found that most experiments can probe regions beyond the current bounds allowed by the EWPT. For $f_{\hat{\pi}}=1$ GeV, the reaches of the dark shower searches are typically somewhat stronger than the FCNC decay searches, although for larger $f_{\hat{\pi}}$ the trend could be reversed. CMS scouting search and LHCb generally give comparable limits, with LHCb winning in the case of a light $Z^{\prime}$. For longer decay lengths, the auxiliary detectors, in particular, MATHUSLA, will have the most powerful reach thanks to its location and large volume. Of course, these two benchmark models only represent a small set of the large possible model space, and the comparison of different experiments within these models should not be taken as general conclusions. Nevertheless, they demonstrate how the results obtained in this paper can be applied to specific models.
@@ Appendix A Dimuon Vertex Trigger Efficiencies at HL-LHC
Appendix B Benchmarks for Model-Specific Projections